Investors require a 17% rate of return on Levine Company's stock (i.e., rs = 17%).

  1. What is its value if the previous dividend was D0 = $2.50 and investors expect dividends to grow at a constant annual rate of (1) -4%, (2) 0%, (3) 5%, or (4) 11%? Do not round intermediate calculations. Round your answers to two decimal places.

    (1) $

    (2) $

    (3) $

    (4) $

please help asap!

Thank you!
Answer #1

Value of stock=D0*(1+g)/(r-g)

1.
=2.5*(1-4%)/(17%-(-4%))
=11.42857143

2.
=2.5*(1+0%)/(17%-0%)
=14.70588235

3.
=2.5*(1+5%)/(17%-5%)
=21.875

4.
=2.5*(1+11%)/(17%-11%)
=46.25